OpenStudy (anonymous):

is y=x-9 an inversed function

OpenStudy (ybarrap):

To check, translate y into x and x into y, like this $$ x=y-9 $$ Now solve for y: $$ y=x+9 $$ If it is an inverse function then taking this function plugging in y=x-9 for x should get you back to x: $$ y=x+9\\ y=(x-9)+9\\ =x $$ Since the inverse function of y=x-9 gave us back x, then it is an inverse function of y=x+9. Notice that we can do the reverse. Plug in y=x+9 into your original equation and you should get back x: $$ y=x-9\\ =(x+9)-9\\ =x $$ Which shows that y=x-9 is the inverse function of y=x+9. Essentially ones takes you from the x domain the to y domain. The other takes you back to the x domain and gives you back what you started with, namely x. Hope this helps.

OpenStudy (anonymous):

It did thanks